OpenStudy (mathmath333):

Their are 4 letters and 4 envelopes.find the number of ways in which only two letters go in wrong envelope ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{Their are 4 letters and 4 envelopes.find the number of ways }\hspace{.33em}\\~\\ & \normalsize \text{in which only two letters go in wrong envelope ?}\hspace{.33em}\\~\\ & a.)\ 4 \hspace{.33em}\\~\\ & b.)\ 5 \hspace{.33em}\\~\\ & c.)\ 6 \hspace{.33em}\\~\\ & d.)\ 3 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

its c 6 ways

OpenStudy (mathmath333):

yea how u got it

OpenStudy (anonymous):

we will need to select the two letters which have to be put in wrong envelop . that is \[4 \times \frac{ 3 }{ 2 }\] i.e \[^{4}c _{2} = 6\]

OpenStudy (mathmath333):

thnks

OpenStudy (anonymous):

no worries :)

OpenStudy (anonymous):

Yep :) It's C, 6 ways :)